7 people go to cinema by their own bike and input their keys in a bag after completation of cinema 
Only three of them start their bike .HOW many 
Ways this can be done .

Grade:12th pass

1 Answers

Arun
3 years ago
Dear student
 
no of ways of this happening = 7C3 *(De arrangement of rest 4 people)= 7C3* 9 = 35 *9 = 315
 
Hope it helps
 
Regards
Arun (askIITians forum expert)